Ask
your own question, for FREE!
Mathematics
16 Online
A television game has 6 shows doors, of which the contests must pick 2.behind two of the doors are expensive cars, and behind the other 4 doors are consolation prizes. Find the probl that the contestant wins exactly 1 car, no car, or atleast one car
Still Need Help?
Join the QuestionCove community and study together with friends!
Find the prob the contest wins exacly one car
there are 6 choose 2 ways to pick 2 doors. 6 choose 2 is 6*5/2 = 15. this is your denominator. there are 8 ways to pick one car and one consolation prize. answer to the first question is 8/15 there are 4 choose 2 ways to pick 2 out of the 4 consolation prizes. 4 choose 2 is 6. answer to the second question is 6/15. at least one car is the same as 'not zero cars' so that probability is 1 - 6/15 = 9/15 = 3/5
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
13 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal